OpenStudy (anonymous):

Evaluate the integral 60e^(3.4x) (30,50)

OpenStudy (lgbasallote):

\[\int 60e^{3.4x}dx \implies 60 \int e^{3.4x}dx\]

OpenStudy (anonymous):

then what would be the next step?

OpenStudy (lgbasallote):

then you integrate e^3.4x

OpenStudy (anonymous):

its supposed to be 6e^(3.4x) btw my mistake

OpenStudy (anonymous):

but the integral of that would be (5e*(17x/5))/17 then?

OpenStudy (anonymous):

(5e^(17x/5))/17

OpenStudy (lgbasallote):

yup..now multiply that by 30

OpenStudy (lgbasallote):

i mean60

OpenStudy (anonymous):

i meant 6 so 6 times that would be 300.1115556

OpenStudy (anonymous):

but then where would the (30,50) come into play?

OpenStudy (lgbasallote):

hmm where did you get that?

OpenStudy (anonymous):

they are they constraints

OpenStudy (lgbasallote):

\[60 \times \frac{5e^{17x/5}}{17}\] \[\frac{60 \times 5}{17} \times e^{17x/5}\]

OpenStudy (anonymous):

\[\int\limits_{30}^{50}\]

OpenStudy (lgbasallote):

did you plug in the limits already?

OpenStudy (anonymous):

no i havent

OpenStudy (anonymous):

im not sure how

OpenStudy (lgbasallote):

do \[\frac{60 \times 5}{17} \times e^{17x/5}\] first

OpenStudy (lgbasallote):

what is 60 x 5/17?

OpenStudy (anonymous):

its supposed to be 6 lol that was my mistake from the beginning in my first post so its supposed to be 6* 5/17

OpenStudy (lgbasallote):

ohhh

OpenStudy (anonymous):

which is 1.764705882

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

\[1.77 e^{17x/5}\] now plug in the limits

OpenStudy (anonymous):

how would i plug in the limits?

OpenStudy (lgbasallote):

\[\large 1.77[e^{(17 \times 50)/5} - e^{(17 \times 30)/5}]\]

OpenStudy (lgbasallote):

i have to go now sorry...hope you figure out what that means :D

OpenStudy (lgbasallote):

but basically i just did \[\large [f(x)|_a^b \implies f(b) - f(a)\]

OpenStudy (lgbasallote):

got that?

OpenStudy (anonymous):

yeahh just coming up a huge number but i think i understand thanks

OpenStudy (lgbasallote):

you're welcome! :)